OpenStudy (anonymous):

i don't understand this sigma question? i have the answer (description) PLEASE HELP SIGMA!!

OpenStudy (anonymous):

\[\sum_{4}^{16} t_{k} = 163 and \sum_{4}^{16} u_{k} = -48, evaluate \sum_{4}^{16} (3+2t _{k} -5u _{k})\]

OpenStudy (anonymous):

the next step is \[\sum_{4}^{16} 3 + \sum_{4}^{16} 2t _{k} + \sum_{4}^{16} (-5u _{k})\]

OpenStudy (anonymous):

but idk why

jimthompson5910 (jim_thompson5910):

They used the rule \[ \Large \sum(x+y) = \sum(x) + \sum(y)\]

OpenStudy (anonymous):

where?

jimthompson5910 (jim_thompson5910):

Like this \[\Large \sum_{4}^{16} (3+2t _{k} -5u _{k}) = \sum_{4}^{16}( (3+2t _{k}) + (-5u _{k}))\] \[\Large \sum_{4}^{16} (3+2t _{k} -5u _{k}) = \sum_{4}^{16}(3+2t _{k}) + \sum_{4}^{16}(-5u _{k})\] \[\Large \sum_{4}^{16} (3+2t _{k} -5u _{k}) = \sum_{4}^{16}(3)+\sum_{4}^{16}(2t _{k}) + \sum_{4}^{16}(-5u _{k})\]

OpenStudy (anonymous):

oh wow, thanks for clearing it up!

jimthompson5910 (jim_thompson5910):

you're welcome